OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <10, 6>, v = <9, 5>

OpenStudy (anonymous):

Use the dot product: \[u \cdot v = |u||v|\cos \theta\] If they are parallel, \(\theta = 0\) or \(\theta = \pi\), and then \[u \cdot v = |u||v|\cos 0 = |u||v|\] or \[u \cdot v = |u||v|\cos \pi = -|u||v|\] If they are orthogonal, \(\large \theta = \frac{\pi}{2}\) and \[u \cdot v = |u||v|\cos \frac{\pi}{2} = 0\] So if \(u \cdot v\) equals \(\pm|u||v|\) they are parallel, and if \(u \cdot v = 0\) they are parallel. If it's not one of those they are neither

OpenStudy (anonymous):

thankyou very much! would you mind helping me on one more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (anonymous):

Again you can use the dot product, but this time we'll use it a bit differently \[u \cdot v = |u||v|\cos \theta\] Now we want to find the angle \(\theta\), so we isolate \(\cos \theta\) \[\cos \theta = \frac{u \cdot v}{|u||v|}\] The angle is simply the inverse cosine \[\theta = \cos^{-1}\left(\frac{u \cdot v}{|u||v|}\right)\]

OpenStudy (anonymous):

so how do i plug the given coordinates in?

OpenStudy (anonymous):

The dot product for two two-dimensional vectors can be calculated by using \[<u_1,u_2> \cdot <v_1, v_2> = u_1v_1+u_2v_2\] And |u| and |v| mean the magnitude of the vectors, which can be calculated by using \[|<u_1, u_2>| = \sqrt{u_1^2 + u_2^2} \] notice that this is basically just the Pythagorean theorem :p

OpenStudy (anonymous):

3.0° 6.0° -7.0° 16.0 are the only answers given. i really still don't understand it im sorry

OpenStudy (anonymous):

We want to know the angle between the 2 vectors u = <2, -4> and v = <3, -8> The first step is calculating the dot product using the method I listed above u \(\cdot\) v = <2, -4> \(\cdot\) <3, -8> = (2 * 3) + (-4 * -8) = 38 The dot product is also equal to the product of the magnitudes, times the cosine of the angle between the two vectors (u \(\cdot\) v = |u||v|cos \(\theta\)) u \(\cdot\) v = \(\sqrt{2^2 + (-4)^2}\sqrt{(3)^2 + (-8)^2}\cos \theta = \sqrt{20}\sqrt{73}\cos \theta\) So setting those equal we get \[38 = \sqrt{20}\sqrt{73} \cos \theta\] \[\cos \theta = \frac{38}{\sqrt{20}\sqrt{73}}\] \[\theta = \cos^{-1} \frac{38}{\sqrt{20}\sqrt{73}}\] Which is around 6 degrees. Sorry for the late reply by the way!

OpenStudy (anonymous):

THANKYOU THANKYOU THANKYOU you're a life saver :)